1) If a star has an apparent magnitude of -0.4 and a parallax of 0.3”, what is a) the distance modulus and b) the absolute magnitude?
Answer #1

Distance Modulus is given by

m -M = 5 log10d - 5

here m = Apparent magnitude = - 0.4

M = Absolute magnitude = ?

Parallax = 0.3"

Relation between distance d and Parallax is given by d = 1 / p

d = 1 / 0.3 = 3.33 parsec

a) Distance Modulus = 5 log10(3.33) - 5

= 2.612 - 5

= - 2.388

b) The absolute magnitude M = m - Distunce modulus

= - 0.4 - ( - 2.388)   

= - 0.4 + 2.388

= 1.988